OpenStudy (anonymous):

................................................................................................................................ What is 417 divided by 21 times 6...........

OpenStudy (chantysquirrel1129**):

You can use a calculator lol :D

OpenStudy (bela):

119.14

OpenStudy (alexandervonhumboldt2):

use your brain

OpenStudy (chantysquirrel1129**):

No direct answers pls ._.

OpenStudy (chantysquirrel1129**):

hahaha Shurik XD

OpenStudy (smilie2015):

119.14285714

OpenStudy (bela):

lol @AlexandervonHumboldt2

jabez177 (jabez177):

She just wanted a medal... xD Welcome to OpenStudy, @Bela! 1 of our rules here is we're supposed to guide the person to the answer, not just give it. :)

OpenStudy (bela):

I don't want a medal @jabez177

OpenStudy (anonymous):

crazy

OpenStudy (alexandervonhumboldt2):

1/2smart+1/2crazy=????

OpenStudy (alexandervonhumboldt2):

1/2*(smart+crazy)=mytery

OpenStudy (anonymous):

u + ur q = habal